# Reconstructive Visual Instruction Tuning

### v1

@@ 2 Related Work
Visual Instruction Tuning. Most visual instruction tuning-based LMMs adopt a plug-in architecture (Liu et al., 2023a; 2024a; Bai et al., 2023), where a language-supervised visual encoder (Radford et al., 2021; Zhai et al., 2023) is responsible for extracting visual tokens. A connector is used to map those visual representations into the LLM space, e.g., Resamplers (Alayrac et al., 2022), Q-Formers (Li et al., 2023b; Dai et al., 2023; Bai et al., 2023; Ge et al., 2024a), and MLPs (Liu et al., 2023a; 2024a; Li et al., 2024c; Liu et al., 2024b; Li et al., 2024a). These LMMs usually follow a two-stage training recipe. During the alignment stage, the connector is trained on high-quality caption data. Next, the full model is trained on single-image visual instruction tuning data. However, only text outputs are supervised. Ross, on the other hand, introduces novel vision-centric supervision via reconstructing fine-grained visual tokens conditioned on visual outputs.
@@ 3 Preliminaries
where is the parameters and indicates the projected visual tokens. is the number of visual tokens and indicates the feature channel. The visual encoder could be either frozen (Liu et al., 2023a; 2024a; Tong et al., 2024a) or fine-tuned (Liu et al., 2024b; Bai et al., 2023; Li et al., 2024c; Wang et al., 2024b).
@@ 5.1 Ablation Study
Finally, we leverage the insights and conclusions from all our previous studies to train our Ross. Specifically, we regard the optimal formulation Ross (KL-16), i.e., denoising with the KL-16 (Rombach et al., 2022) tokenizer, as our final Ross.
@@ 5.2 In-Depth Analysis
Generative v.s. Reconstructive. We ablate the effectiveness of reconstruction over generation in Table 3. Similar to Sun et al. (2024b) and Dong et al. (2024), for generative cases, we adopt 576 learnable latent tokens to query the LMM and utilize the corresponding outputs as conditions to the denoiser. The detailed pipeline of these two methods can be found at Figure 11 in Appendix C. However, generative methods require specific creation data and can not naively implemented on the original SFT data, i.e., Cambrian-737K (Tong et al., 2024a). To build creation data, we utilize GPT-4o to transfer 102K caption into text-to-image creation data from ShareGPT4V (Chen et al., 2023b) and combine them with the original SFT data. From Table 3, we can tell that reconstructive objectives boost comprehension while generative alternatives cannot.
Ross with Different LLMs and Visual Encoders. To demonstrate the effectiveness of our vision-centric supervision adopted by our Ross, we conduct systematic experiments across different base LLMs and visual encoders. From Table 3, Ross contributes to significant improvements consistently, especially on fine-grained comprehension benchmarks, i.e., MMVP (Tong et al., 2024b) and ChartQA (Masry et al., 2022).
@@
| Ross-7B | 89.2 | 57.3 | 79.0 | 76.1 | 73.0 | 43.4 | 54.7 | 65.5 | 76.8 |
@@ 5.3 Comparison with State-of-the-Arts
Ross utlizes a single SigLIP-ViT-SO400M/14@384 (Zhai et al., 2023) as the visual encoder. Ross-7B utilizes Qwen2-7B-Instruct (Yang et al., 2024a) and Ross-13B adopts Vicuna-13B-v1.5 (Chiang et al., 2023) as the base LLM. The implementation almost follows LLaVA-v1.5 (Liu et al., 2024a) without the high-resolution image-slicing technique (Liu et al., 2024b). Thus, our primary comparison of Ross with alternative methods focuses on benchmarks that do not require exceptionally high-resolution inputs. When comparing with other methods, we incorporate more training data. Specifically, we use a mixture of 2M caption data for the pre-training stage, which consists of 1246K from ShareGPT4V (Chen et al., 2023b) and 707K from ALLaVA (Chen et al., 2024a). The instruction tuning data is a mixture of Cambrian-737K (Tong et al., 2024a) and SMR-473K (Zhang et al., 2024).
@@
5.4 Transfer Learning on Understanding Depth Maps
We further evaluate the transfer learning capability of our Ross on SpatialBench (Cai et al., 2024), which requires the model to understand depth maps. We compare our Ross with the visual instruction tuning baseline, with the same training data and model architecture. Also, we compare the effectiveness of the extrinsic assistance solution, i.e., combining a depth expert MiDaS-3.0 (Birkl et al., 2023) to visual instruction tuning, with our intrinsic activation solution. Specifically, the pre-training data is LLaVA-558K (Liu et al., 2023a) and the fine-tuning data is SpatialQA-853K (Cai et al., 2024), where each conversation contains the RGB image and the depth maps extracted by ZoeDepth (Bhat et al., 2023). The visual encoder is CLIP-ViT-L/14@336 (Radford et al., 2021) and the base LLM is Qwen2-7B-Instruct (Yang et al., 2024a).
As demonstrated in Table 5, our Ross manages to make use of the extra depth map, as consistent and significant improvements are observed when taking “RGB + D” inputs for testing. Extrinsic assistance approaches cannot take advantage of extra depth maps when testing. Even GPT-4o cannot fully understand depth maps. Figure 9 provides qualitative results. The extra depth understanding visual expert, i.e., MiDaS (Birkl et al., 2023), fails to help LMMs understand depth maps both quantitatively in Table 5 and qualitatively in Figure 9.
@@ 6 Conclusion
This paper introduces reconstructive visual instruction tuning (Ross). Different from conventional LMMs that solely utilize text supervision, Ross leverages a vision-centric reconstructive objective to supervise visual outputs. To avoid being overwhelmed by heavily redundant raw RGB values, we train a denoiser to recover clean latent visual representations conditioning on visual outputs. Experimentally, the proposed objective indeed brings enhanced comprehension capabilities and reduced hallucinations. Ross outperforms the state-of-the-art under most cases with only a single SigLIP (Zhai et al., 2023) as the visual encoder. In-depth analysis demonstrates the effectiveness of Ross for directing focus towards visual elements and understanding depth maps.
Discussion. One limitation is that Ross does not have generation capabilities, since Ross is designed for enhanced multimodal comprehension, without the need to generate photorealistic aesthetic images. Furthermore, the gap in training data between comprehension and generation methods also matters. For instance, PixArt- (Chen et al., 2023a), which is one of the most efficient text-to-image models, was trained on nearly 400M images to model the pixel discrepancy just in the first training stage. By contrast, our Ross is only trained on nearly 3M images for one epoch. Future topics include achieving photorealistic text-to-image generation via incorporating more training samples.

### v2

@@ 2 Related Work
Visual Instruction Tuning. Most visual instruction tuning-based LMMs adopt a plug-in architecture (Liu et al., 2023a; 2024a; Bai et al., 2023; Wang et al., 2024a), where a language-supervised visual encoder (Radford et al., 2021; Zhai et al., 2023) is responsible for extracting visual tokens. A connector is used to map those visual representations into the LLM space, e.g., Resamplers (Alayrac et al., 2022), Q-Formers (Li et al., 2023b; Dai et al., 2023; Bai et al., 2023; Ge et al., 2024a), and MLPs (Liu et al., 2023a; 2024a; Li et al., 2024c; Liu et al., 2024b; Li et al., 2024a). These LMMs usually follow a two-stage training recipe. During the alignment stage, the connector is trained on high-quality caption data. Next, the full model is trained on single-image visual instruction tuning data. However, only text outputs are supervised. Ross, on the other hand, introduces novel vision-centric supervision via reconstructing fine-grained visual tokens conditioned on visual outputs.
@@ 3 Preliminaries
where is the parameters and indicates the projected visual tokens. is the number of visual tokens and indicates the feature channel. The visual encoder could be either frozen (Liu et al., 2023a; 2024a; Tong et al., 2024a) or fine-tuned (Liu et al., 2024b; Bai et al., 2023; Li et al., 2024c; Wang et al., 2024c).
@@ 5.1 Ablation Study
Finally, we leverage the insights and conclusions from all our previous studies to train our Ross. Specifically, we regard the optimal formulation Ross (KL-16), i.e., denoising with the KL-16 tokenizer, as our final Ross. Please refer to Section C.1 for ablations on the architecture of the denoiser, continuous tokenizer v.s. discrete tokenizer, and the denoising schedule.
@@ 5.2 In-Depth Analysis
Generative v.s. Reconstructive. We ablate the effectiveness of reconstruction over generation in Table 3. Similar to Sun et al. (2024b) and Dong et al. (2024), for generative cases, we adopt 576 learnable latent tokens to query the LMM and utilize the corresponding outputs as conditions to the denoiser. The detailed pipeline of these two methods can be found at Figure 11 in Section C.1. However, generative methods require specific creation data and can not be naively implemented on the original SFT data. To build creation data, we utilize GPT-4o to transfer 102K caption into text-to-image creation data from ShareGPT4V (Chen et al., 2023b) and combine them with the original SFT data. From Table 3, we can tell that reconstructive objectives boost comprehension while generative alternatives cannot. An intuitive explanation of this evidence can be found in Section C.1.
Ross with Different LLMs and Visual Encoders. To demonstrate the effectiveness of our vision-centric supervision adopted by our Ross, we conduct systematic experiments across different base LLMs and visual encoders. From Table 3, Ross contributes to significant improvements consistently, especially on fine-grained comprehension benchmarks, i.e., MMVP (Tong et al., 2024b) and ChartQA (Masry et al., 2022). Extended experiments on more representative benchmarks can be found at Table 12 in Section C.1.
@@
| Ross-7B | 88.3 | 57.1 | 79.1 | 77.1 | 73.6 | 46.6 | 56.7 | 65.5 | 79.3 |
| Base LLM: Vicuna-7B-v1.5 | |||||||||
| LLaVA-v1.6-7B‡ (Liu et al., 2024b) | 86.5 | 35.8 | 67.4 | 60.1 | 70.2 | 35.8 | 37.3 | 64.2 | 67.1 |
| Ross-7B | 88.2 | 55.2 | 67.7 | 61.3 | 67.6 | 36.9 | 39.3 | 63.7 | 69.3 |
@@
Reconstruction Results. We fine-tune the denoiser to recover latent tokens from a frozen KL-16 provided by Rombach et al. (2022) conditioned on Ross-7B features on ImageNet-1K (Deng et al., 2009) for only five epochs, where the denoiser manages to produce reasonable reconstruction results as illustrated in Figure 9. This interesting finding demonstrates that high-level Ross-7B features actually contain image details. We hope this finding will inspire future work.
Computational Overhead. The denoising process introduces a negligible increase in training time (10% compared to the baseline), while the benefits outweigh the minor additional costs. Please refer to Table 10 in Appendix B for details.
@@ 5.3 Comparison with State-of-the-Arts
Ross utlizes a single SigLIP-ViT-SO400M/14@384 (Zhai et al., 2023) as the visual encoder. Ross-7B utilizes Qwen2-7B-Instruct (Yang et al., 2024a) and Ross-13B adopts Vicuna-13B-v1.5 (Chiang et al., 2023) as the base LLM. The implementation almost follows LLaVA-v1.5 (Liu et al., 2024a) without the high-resolution image-slicing technique (Liu et al., 2024b). Thus, our primary comparison of Ross with alternative methods focuses on benchmarks that do not require exceptionally high-resolution inputs. We use a mixture of 2M caption data for the pre-training stage, which consists of 1246K from ShareGPT4V (Chen et al., 2023b) and 707K from ALLaVA (Chen et al., 2024a). The instruction tuning data is a mixture of Cambrian-737K (Tong et al., 2024a) and SMR-473K (Zhang et al., 2024). We further incorporate our Ross with the “anyres” technique (Liu et al., 2024b) and compare with others on high-resolution benchmarks at Table 13 in Section C.2.
@@
5.4 Applications
Transfer Learning on Understanding Depth Maps. We further evaluate the transfer learning capability of our Ross on SpatialBench (Cai et al., 2024), which requires the model to understand depth maps. We compare our Ross with the visual instruction tuning baseline, with the same training data and model architecture. Also, we compare the effectiveness of the extrinsic assistance solution, i.e., combining a depth expert MiDaS-3.0 (Birkl et al., 2023) to visual instruction tuning, with our intrinsic activation solution. Specifically, the pre-training data is LLaVA-558K (Liu et al., 2023a) and the fine-tuning data is SpatialQA-853K (Cai et al., 2024), where each conversation contains the RGB image and the depth maps extracted by ZoeDepth (Bhat et al., 2023). The visual encoder is CLIP-ViT-L/14@336 (Radford et al., 2021) and the base LLM is Qwen2-7B-Instruct (Yang et al., 2024a). As demonstrated in Table 5, our Ross manages to make use of the extra depth map, as consistent and significant improvements are observed when taking “RGB + D” inputs for testing. Extrinsic assistance approaches cannot take advantage of extra depth maps when testing. Even GPT-4o cannot fully understand depth maps. Qualitative results can be found at Figure 16 in Section C.4.
@@ 6 Conclusion
This paper introduces reconstructive visual instruction tuning (Ross), leveraging a vision-centric reconstructive objective to supervise visual outputs. To avoid being overwhelmed by heavily redundant raw RGB values, we train a denoiser to recover clean latent visual representations conditioning on visual outputs. Experimentally, the proposed objective indeed brings enhanced comprehension capabilities and reduced hallucinations. Ross outperforms the state-of-the-art under most cases with only a single SigLIP (Zhai et al., 2023) as the visual encoder. The in-depth analysis demonstrates that high-level features from Ross-7B actually contain sufficient details for low-level image reconstruction. This finding reveals the possibility of equipping comprehension LMMs with the ability of naive generation without the help of generation experts such as Stable Diffusion (Rombach et al., 2022).
@@ Appendix B Implementation Details
| Vision | Base LLM |
Speed (s/iter) | Time |
| CLIP-L/336 | Qwen2-7B-Instruct | – | 7.63 B | 8.31 | 13h 17min | 45.34 G | ||||
| CLIP-L/336 | Qwen2-7B-Instruct | ✓ | 7.68 B | 9.02 (1.09) | 14h 25min | 46.62 G (1.03) | ||||
| CLIP-L/336 | Vicuna-13B-v1.5 | – | 13.05 B | 13.33 | 21h 19min | 48.62 G | ||||
| CLIP-L/336 | Vicuna-13B-v1.5 | ✓ | 13.11 B | 14.69 (1.10) | 23h 30min | 49.07 G (1.01) | ||||
| SigLIP-L/384 | Qwen2-7B-Instruct | – | 7.63 B | 8.77 | 14h 1min | 47.08 G | ||||
| SigLIP-L/384 | Qwen2-7B-Instruct | ✓ | 7.68 B | 9.48 (1.08) | 15h 9min | 52.07 G (1.11) | ||||
| SigLIP-L/384 | Vicuna-13B-v1.5 | – | 13.05 B | 14.22 | 22h 44min | 48.80 G | ||||
| SigLIP-L/384 | Vicuna-13B-v1.5 | ✓ | 13.11 B | 15.32 (1.08) | 24h 30min | 52.68 G (1.08) |
@@
Computational Costs. As demonstrated in Table 10, the denoising process introduces a negligible increase in training time (10% compared to the baseline), while the benefits outweigh the minor additional costs.
@@ C.1 More Ablations
| Schedule of | POPE | Hallu. | MMVP | ChartQA | MMB |
| – | 87.9 | 55.0 | 29.3 | 34.0 | 73.8 |
| Linear (Ho et al., 2020) | 88.1 0.2 | 57.3 2.3 | 42.0 12.4 | 39.2 5.2 | 75.1 1.3 |
| Scaled Linear (Rombach et al., 2022) | 88.4 0.5 | 58.3 3.3 | 40.0 10.4 | 40.7 6.7 | 75.3 1.5 |
| GLIDE Softmax (Nichol et al., 2022) | 88.4 0.5 | 59.1 4.1 | 42.7 13.4 | 40.4 6.4 | 75.2 1.4 |
| GeoDiff Sigmoid (Xu et al., 2022) | 88.2 0.3 | 57.7 2.7 | 41.3 11.7 | 38.9 4.9 | 75.5 1.7 |
Schedule of . We study the effectiveness of different schedules of in Table 11. From the table, we can tell that even with different schedules of , Ross consistently improves the baseline, demonstrating its robustness to the denoising schedule.
@@
| CLIP-ViT-L/14@336 | SigLIP-ViT-SO400M/14@384 | |||||||
| Vicuna-7B-v1.5 | Qwen2-7B-Instruct | Vicuna-7B-v1.5 | Qwen2-7B-Instruct | |||||
| Benchmark | LLaVA | Ross | LLaVA | Ross | LLaVA | Ross | LLaVA | Ross |
| POPE | 86.3 | 87.2 0.9 | 87.9 | 88.4 0.5 | 86.0 | 87.7 1.7 | 88.5 | 88.7 0.2 |
| HallusionBench | 52.5 | 55.8 3.3 | 55.0 | 59.1 4.1 | 50.4 | 53.8 3.4 | 57.3 | 58.2 0.9 |
| MMBench-EN | 67.0 | 67.6 0.6 | 73.8 | 75.2 1.4 | 64.5 | 69.2 4.7 | 76.3 | 76.9 0.6 |
| MMBench-CN | 60.0 | 59.8 0.2 | 72.9 | 73.7 0.8 | 63.1 | 63.4 0.3 | 75.7 | 76.3 0.7 |
| SEED | 66.7 | 66.4 0.3 | 70.3 | 70.7 0.4 | 68.2 | 69.0 0.8 | 72.3 | 72.1 0.2 |
| MMMU | 30.0 | 34.0 4.0 | 44.0 | 45.3 1.3 | 33.3 | 38.0 4.7 | 38.7 | 41.3 2.6 |
| MMMU | 35.3 | 36.0 0.7 | 41.9 | 42.6 0.7 | 34.2 | 35.4 1.2 | 41.8 | 43.8 2.0 |
| MMVP | 28.0 | 36.0 8.0 | 29.3 | 42.7 13.4 | 27.3 | 38.0 10.7 | 40.7 | 49.3 8.6 |
| AI2D | 61.2 | 61.4 0.2 | 71.9 | 73.3 1.4 | 62.6 | 62.4 0.2 | 74.0 | 74.5 0.5 |
| ChartQA | 32.9 | 39.8 6.9 | 36.2 | 41.6 5.4 | 34.0 | 48.2 14.2 | 44.4 | 46.9 2.5 |
| DocVQA | 33.4 | 41.6 8.2 | 31.1 | 44.7 13.6 | 40.4 | 40.7 0.3 | 39.2 | 39.3 0.1 |
| InfoVQA | 21.2 | 26.4 5.2 | 22.1 | 39.3 16.2 | 22.8 | 23.3 0.5 | 24.0 | 25.1 1.1 |
| TextVQA | 55.7 | 58.7 3.0 | 52.0 | 54.1 2.1 | 60.5 | 62.6 2.1 | 56.3 | 57.5 1.2 |
| OCRBench | 339 | 350 11 | 363 | 381 18 | 354 | 365 11 | 432 | 448 16 |
| RealWorldQA | 52.7 | 53.2 0.5 | 56.7 | 57.4 0.7 | 55.0 | 57.1 2.1 | 57.9 | 59.1 1.2 |
| Average | 47.8 | 50.6 2.8 | 52.1 | 56.4 4.3 | 49.2 | 52.4 3.2 | 55.4 | 56.9 1.5 |
We hypothesize that the underlying reason for the lower performance of generative methods in comprehension tasks is the weak correspondence between inputs and supervision under generative settings, which typically arises from both the (1) data and the (2) design of these methods.
(1) Typical generative methods that explore the synergy of comprehension and generation, usually leverage image generation conditioned on text instructions on (i) text-to-image datasets or (ii) interleaved datasets as extra supervision. However, (i) text-to-image datasets are typically designed to generate high-aesthetic samples rather than text-aligned ones, and (ii) interleaved datasets aim to enable few-shot learning via interleaving independent supervised examples, where reasoning becomes more important than alignment. Therefore, there exists a clear disconnect where the supervision (image) has little to do with the input (text instruction). For example, the CLIP-Score (Hessel et al., 2021), which measures the similarity between text and images, is only 0.3043 for the LAION-Art dataset (Schuhmann et al., 2022) and 0.2842 for the MMC4 dataset (Zhu et al., 2023), indicating that the supervision signals in these datasets are not well-suited for tasks requiring strong text-image alignment.
(2) Even when we attempt to ensure image-text alignment by converting aligned caption data into creation data for supervision, the results demonstrated in Table 3 remain unsatisfactory. This suggests that the design of generative objectives itself does not inherently require a strong correspondence between inputs and supervision targets.
In contrast, reconstructive methods like Ross leverage the original input images as auxiliary supervision, ensuring a strong and direct correspondence, which is crucial for tasks requiring accurate comprehension and interpretation of multimodal data, leading to significantly improved performance.
Extended Ablations on Different LLMs and Visual Encoders. We extend the ablation in Table 3 by incorporating more benchmarks, providing a more balanced and representative distribution of tasks. Empirical results in Table 12 demonstrate that our proposed vision-centric supervision utilized by Ross leads to significant improvements in most cases. Moreover, we found Ross contributes more significant improvements over fine-grained comprehension datasets, such as HallusionBench (Guan et al., 2024), MMVP (Tong et al., 2024b), ChartQA (Masry et al., 2022), and OCRBench (Liu et al., 2023c).
C.2 Comparison on High-Resolution Benchmarks
We incorporate the “anyres” technique proposed by LLaVA-v1.6 (Liu et al., 2024b) into our Ross. Specifically, for each image, we employ a grid configuration of 384{22, 1{2,3,4}, {2,3,4}1} to identify the input resolution, resulting in a maximum of 5729 3,645 visual tokens. Each 384384 crop is required to reconstruct the original input via the denoising objective proposed by Ross. In Table 13, our Ross-7B surpasses LLaVA-v1.6-7B (Liu et al., 2024b) and Cambrian-1-8B (Tong et al., 2024a) under most cases. These results indicate that Ross not only performs well at lower resolutions but also maintains its competitive edge at higher resolutions, making it a robust and versatile method.
| Model | ChartQA | DocVQA | InfoVQA | TextVQA | OCRBench | RealWorldQA |
| GPT-4V-1106 (OpenAI, 2023a) | 78.5 | 88.4 | – | 78.0 | 645 | 61.4 |
| Gemini-1.5 Pro (Team et al., 2023) | 81.3 | 86.5 | – | 78.1 | – | 67.5 |
| Grok-1.5 (xAI, 2024) | 76.1 | 85.6 | – | 78.1 | – | 68.7 |
| LLaVA-v1.5-7B‡ (Liu et al., 2024a) | 18.2 | 28.1 | 25.7 | 58.2 | 317 | 54.9 |
| LLaVA-v1.6-7B‡ (Liu et al., 2024b) | 65.5 | 74.4 | 37.1 | 64.8 | 532 | 57.6 |
| Cambrian-1-8B (Tong et al., 2024a) | 73.3 | 77.8 | – | 71.7 | 624 | 64.2 |
| Ross-7B | 76.9 | 81.8 | 50.5 | 72.2 | 607 | 66.2 |
| CLIP-ViT-L/14@336 | SigLIP-ViT-SO400M/14@384 | |||||||
| Vicuna-7B-v1.5 | Qwen2-7B-Instruct | Vicuna-7B-v1.5 | Qwen2-7B-Instruct | |||||
| Benchmark | LLaVA | Ross | LLaVA | Ross | LLaVA | Ross | LLaVA | Ross |
| Vision-Language Benchmarks on Knowledge | ||||||||
| ScienceQA | 68.5 | 69.0 0.5 | 76.5 | 77.4 0.9 | 69.6 | 71.3 1.7 | 78.3 | 78.5 0.2 |
| MMMU | 30.0 | 34.0 4.0 | 44.0 | 45.3 1.3 | 33.3 | 38.0 4.7 | 38.7 | 41.3 2.6 |
| MMMU | 35.3 | 36.0 0.7 | 41.9 | 42.6 0.7 | 34.2 | 35.4 1.2 | 41.8 | 43.8 2.0 |
| AI2D | 61.2 | 61.4 0.2 | 71.9 | 73.3 1.4 | 62.6 | 62.4 0.2 | 74.0 | 74.5 0.5 |
| Language Benchmarks | ||||||||
| MMLU | 26.5 | 27.4 0.9 | 57.1 | 60.7 3.6 | 26.0 | 25.9 0.1 | 60.9 | 61.0 0.1 |
| HellaSwag | 27.0 | 26.9 0.1 | 46.4 | 46.2 0.2 | 27.1 | 27.0 0.1 | 45.5 | 46.6 1.1 |
| IFEval | 41.2 | 44.6 3.4 | 47.1 | 49.2 2.1 | 43.6 | 43.8 0.2 | 47.8 | 48.1 0.3 |
| IFEval | 28.7 | 35.3 6.7 | 35.1 | 37.0 1.9 | 32.5 | 33.1 0.6 | 35.3 | 36.2 0.9 |
| Average | 39.8 | 41.8 2.0 | 52.5 | 54.0 1.5 | 41.1 | 42.1 1.0 | 52.8 | 53.8 1.0 |
| Benchmark | 0.5B | 1.5B | 3B | 7B | ||||
| LLaVA | Ross | LLaVA | Ross | LLaVA | Ross | LLaVA | Ross | |
| POPE | 50.0 | 60.4 10.4 | 85.3 | 87.9 2.4 | 87.3 | 88.1 0.8 | 87.9 | 88.4 0.5 |
| HallusionBench | 45.8 | 48.0 2.2 | 48.7 | 49.6 0.9 | 52.2 | 52.2 – 0.0 | 48.7 | 53.7 5.0 |
| MMBench-EN | 55.2 | 60.4 5.2 | 67.5 | 68.2 1.7 | 70.6 | 71.4 0.8 | 75.0 | 75.7 0.7 |
| MMBench-CN | 45.6 | 48.9 3.3 | 62.4 | 63.9 1.5 | 68.0 | 69.1 1.1 | 73.6 | 73.5 0.1 |
| SEED | 55.8 | 55.6 0.2 | 66.3 | 66.8 0.5 | 68.2 | 68.4 0.2 | 70.6 | 71.0 0.4 |
| OCRBench | 229 | 248 19 | 291 | 298 7 | 313 | 308 5 | 334 | 358 24 |
| MMMU | 35.2 | 36.0 0.8 | 44.7 | 45.0 0.3 | 48.7 | 49.0 0.3 | 48.0 | 48.0 – 0.0 |
| MMMU | 38.0 | 40.3 1.7 | 41.8 | 43.6 1.8 | 41.6 | 42.7 1.1 | 47.3 | 48.0 0.7 |
| AI2D | 45.3 | 46.0 0.7 | 59.0 | 59.5 0.5 | 62.9 | 63.2 0.3 | 68.3 | 68.5 0.2 |
| RealWorldQA | 45.1 | 46.4 1.3 | 50.5 | 53.5 3.0 | 55.7 | 57.9 2.2 | 59.5 | 59.9 0.4 |
| Average | 43.9 | 46.7 2.8 | 55.3 | 56.8 1.5 | 58.9 | 59.3 0.4 | 61.2 | 62.3 1.1 |
| PT | SFT | POPE | Hallu. | ChartQA | OCRBench | MMB | AI2D | |
| 558K | 737K | – | 87.9 | 55.0 | 34.0 | 363 | 73.8 | 72.4 |
| ✓ | 88.4 0.5 | 59.1 4.1 | 40.4 6.4 | 380 17 | 75.2 1.4 | 73.3 0.9 | ||
| 558K | 1.2M | – | 88.5 | 57.3 | 37.0 | 389 | 75.7 | 74.5 |
| ✓ | 88.8 0.3 | 57.8 0.5 | 42.0 5.0 | 392 3 | 76.8 1.1 | 74.7 0.2 | ||
| 2M | 737K | – | 88.1 | 55.6 | 37.3 | 384 | 76.2 | 72.3 |
| ✓ | 88.3 0.2 | 56.2 0.6 | 41.9 4.5 | 398 14 | 77.0 0.8 | 73.4 1.1 | ||
| 2M | 1.2M | – | 88.5 | 53.8 | 41.2 | 388 | 76.5 | 73.9 |
| ✓ | 88.9 0.4 | 57.3 2.5 | 43.2 2.0 | 405 17 | 78.0 1.5 | 74.1 0.2 |
C.3 More Analysis
Language Capabilities. One possible concern of Ross is that this type of low-level reconstruction may harm the high-level language capabilities. To investigate this issue, we evaluate multi-modal benchmarks that mainly require general knowledge following (Tong et al., 2024a), including ScienceQA (Saikh et al., 2022), MMMU (Yue et al., 2024), and AI2D (Hiippala et al., 2021). Furthermore, we incorporate representative language benchmarks, including general understanding on MMLU (Hendrycks et al., 2020) and HellaSwag (Zellers et al., 2019), and instruction-following on IFEval (Zhou et al., 2023). Empirical results in Table 14 demonstrate that Ross does not harm language capabilities as it brings improvements in most cases.
Model Scaling Properties. To study the stability and scalability of Ross across different model sizes, we use the Qwen2.5 series (Team, 2024) with varying sizes as the base language model while keeping the CLIP-ViT-L/14@336 (Radford et al., 2021) as the visual encoder. The pre-training data is LLaVA-558K (Liu et al., 2023a), and the instruction tuning data is LLaVA-665K (Liu et al., 2024a). The results, shown in Table 16, demonstrate that Ross brings improvements over the baseline (LLaVA) across different model sizes in most cases.
Data Scaling Properties. To study the impact of the training data scale, we used Qwen2-7B-Instruct (Yang et al., 2024a) as the base language model and CLIP-ViT-L/14@336 (Radford et al., 2021) as the visual encoder. We compared the performance of Ross and the baseline under different scales of training data. Table 16 demonstrates that Ross consistently brings significant improvements as the training data scale increases.
Gradient Analysis. To better explain the reasoning behind how the vison-centric supervision enables the model to focus on relevant areas of the image during VQA tasks, we provide qualitative comparison using GradCAM (Selvaraju et al., 2020) on MMVP (Tong et al., 2024b) in Figure 12, since GradCAM helps in understanding which parts of the image the model is focusing on, making the model’s decision-making process more transparent. In our analysis, we visualize the gradient of the second-to-last block of the LMM, regarding the option of the ground-truth answer as the target class. Specifically in this case, where the providing question is about the spider web, our proposed vision-centric supervision signals provide more reasonable gradients and urge LMMs to focus on relevant regions, i.e., the spider web, as the training evolves.
C.4 Qualitative Comparisons